Luna1991:

A 3 3/4 ounce serving of Norway Sardines in chili sauce provides 100% of the U.S recommended daily allowance for vitamin D. what percent of the U.S. daily allowance is provided per ounce of the Norway Sardines in chili sauce?

Luna1991:

i’d really appreciate ur help😔

jhonyy9:

like a first step change 3 3/4 = 15/4 than us the rule of 3 simple and write 15/4 ounces provides 100% 1 ounce provide x% x = 1× 100/ (15/4) = 26,66% so from this result that 1 ounce provide 26,66% hope this is understandably and helped

Luna1991:

that helped a lot thank u very much!